OpenStudy (eli-ite):

A jar contains nine blue marbles, eight yellow marbles, 11 black marbles, and 12 red marbles. Without looking, what is the probability of pulling a yellow marble from the jar?

OpenStudy (queenie):

first, tell me how many total marbles are there ?

OpenStudy (alivejeremy):

simplify 8/40 = 1/5

OpenStudy (alivejeremy):

Answer is 1/5

OpenStudy (eli-ite):

ok its 40

OpenStudy (alivejeremy):

Answer is 1/5

OpenStudy (queenie):

thats ur denominator......and there is 8 yellow ones...thats ur numerator....8/40 reduces to 1/5. Basically, it is set up like this : yellow marbles / total marbles

OpenStudy (alivejeremy):

8 / 40 = 23 / (23 * 5) = (23 : 23) / ((23 * 5) : 23) = (8 : 8) / (40 : 8) = 1 / 5

OpenStudy (eli-ite):

ooo i get ok thanx

OpenStudy (queenie):

yw

OpenStudy (alivejeremy):

You're Welcome! :D